A jar contains 10 black marbles and 40 white marbles.
a. If you randomly select a marble from the jar, what is the probability that you will get a white marble?
b. If you are selecting a random sample of n = 3 marbles, and the first 2 marbles are both white, what is the probability that the third marble will be black?

Answer by edjones(8007)

a. 40/50=4/5
.
b. 38/48=19/24
.
Ed